Lukistan Balance of Payments
Components of the Balance of Payments

($)

Exports of goods and services

+ 440

Merchandise exports (including military sales)

+ 280

Exports of services

+ 40

Income Received from U.S. assets abroad

+ 120

Imports of goods and services

– 490

Merchandise imports (including military purchases)

– 360

Imports of services

– 60

Income Received from foreign assets in U.S.

– 70

Net unilateral transfers abroad

– 11

Outflow of U.S. capital

– 26

Inflow of foreign capital

+ 80

Increase in U.S. official reserve assets

– 14

Increase in foreign official assets in U.S.

+ 21

2. The Lukistan current account balance equals __________ billions of dollars.

3. The Lukistan capital account balance equals __________ billions of dollars (include changes in official reserves as part of capital flows).

5. The sum of the balance of payments for Lukistan, including both the current account, the capital account, and the statistical discrepancy is:
Answer #1

current account will include:

Exports of goods and services

+ 440

Merchandise exports (including military sales)

+ 280

Exports of services

+ 40

Income Received from U.S. assets abroad

+ 120

Imports of goods and services

– 490

Merchandise imports (including military purchases)

– 360

Imports of services

– 60

Income Received from foreign assets in U.S.

– 70

Net unilateral transfers abroad

– 11

hence the total of the current account will be 880-999 = -119

the capital account will include :

Outflow of U.S. capital

– 26

Inflow of foreign capital

+ 80

Increase in U.S. official reserve assets

– 14

Increase in foreign official assets in U.S.

+ 21

hence the total capital account is 101-40 = 61

sum of BOP including both the current account and the capital account is -119+61 = -58

since BOP aggregate must be equal to 0, hence our statistical discrepancy is -58